# Celestial Symmetries of Black Hole Horizons

### v1

@@ I Introduction
The goal of this Letter is to identify the celestial $Lw_{1+\infty}$ symmetries of black hole horizons. More precisely, we establish two main results: $(i)$ We demonstrate a precise correspondence between the Ashtekar-Streubel phase space at null infinity and the subleading phase space near a null hypersurface, such as a black hole horizon. This is achieved using Penrose’s conformal compactification, the behavior of the Newman-Penrose (NP) formalism under Weyl rescalings, and covariant phase space techniques. $(ii)$ Leveraging this correspondence, we identify the $Lw_{1+\infty}$ charges at the horizon and show that, upon imposing self-duality conditions, their fluxes integrated over the horizon yield the canonical generators of the celestial symmetries. In particular, in the absence of radiation through the horizon, these charges are conserved, giving rise to an infinite tower of observables relevant for black hole physics.
@@ II NP formalism at null infinity
$m=\frac{1}{r_{\mathscr{I}}}m_{0}^{A}\partial_{A}+\mathcal{O}(r_{\mathscr{I}}^{-1}),\quad\bar{m}=\frac{1}{r_{\mathscr{I}}}\bar{m}_{0}^{A}\partial_{A}+\mathcal{O}(r_{\mathscr{I}}^{-1})$
@@
    where $\eth$ encodes angular derivative on the sphere [51]. Each $Q_{s}$ for $s>-2$ corresponds to a charge aspect of spin weight $s$ and $Q_{-2}\equiv\Psi^{0}_{4}=\partial_{v}\lambda_{1}$ is the radiation responsible for the non-conservation. These recursion relations correspond to an infinite tower of flux-balance laws at null infinity. For $s=-2,\ldots,2$, we have $Q_{s}=\Psi^{0}_{2-s}$ and Equation (5) follows from the Bianchi identities. For $s>2$, the $Q_{s}$’s appear in the $1/r_{\mathscr{I}}$ expansion of $\Psi_{0}$. Self-dual gravity and full gravity exhibit the same pattern only at leading orders near null infinity. In [26], deviations with respect to the self-dual sector were observed for $s\geq 4$, and obtaining (5) from the full gravity theory requires a truncation of the Bianchi identities.
@@
[…]
@@ III From null infinity to the horizon
Acting with the bare derivative operators $\partial_{v}$ and $\eth$ on weighted scalars does not produce well-defined Weyl weighted scalars. To remedy this issue, one can introduce the conformal GHP derivative operators [58, 57]:
@@
constructed from the bare operators by correcting them with spin coefficients, playing the role of a Weyl connection, so that they have well-defined weights.
We now have all the tools in hand to map the relevant part of the solution space at null infinity to the one at the horizon. We implement the Penrose conformal compactification [53, 54, 55, 56] on the off-shell radial expansions presented in Section II, as the Einstein equations themselves are not Weyl covariant. Starting from the radial expansion of the tetrad (2), and performing the compactification with conformal factor $\Omega$, see (7), we find
@@
$\sigma\to\sigma^{\prime}_{0}+\mathcal{O}(r),\quad\lambda\to\lambda^{\prime}=\lambda_{0}+r\lambda_{1}+\mathcal{O}(r^{2}).$ (12)
@@
Using the conformal GHP operators (9), one can “covariantize” the recursion relations (5) to obtain
þ${\text{\th}}^{\prime}_{\mathscr{C}}Q_{s}=\eth_{\mathscr{C}}Q_{s-1}-(s+1)\sigma_{0}\,Q_{s-2}$ (13)
This covariantization is harmless at null infinity, as all the correction terms involved in this procedure vanish at $\mathscr{I}$. Exploiting the Weyl covariance of these expressions, we can perform the conformal rescaling and re-interpret these equations at the horizon. However, at finite distance, the correction terms in the GHP derivative operators with respect to the bare operators are crucial as the intrinsic geometry at the horizon is non-trivial. For example, one of these corrections involves the surface gravity of the black hole. Hence, (13) constitutes the recursion relations at the horizon and, in particular, $Q_{s}=\Psi_{2-s}^{0}$, $s=-2,\ldots,2$ are now the components of the Weyl tensor in the Taylor expansion (11), and the $Q_{s}$’s for $s>2$ appear in the expansion of $\Psi_{0}$.
@@ IV Subleading phase space
As mentioned in Section II, a key ingredient to identify the $Lw_{1+\infty}$ symmetries from a spacetime perspective at null infinity are self-duality conditions. These are necessary only at subleading order, where the deviation between self-dual and full gravity starts to occur [25, 26]. At finite distance, self-duality conditions are already visible at leading order in $r$ since the intrinsic geometry on the null hypersurface contains genuine degrees of freedom involving both helicities. We propose conditions on the spin coefficients:
@@
where $\mu_{0}$ is the expansion of $\mathcal{H}$ and $\text{Re}(\gamma_{0})=-2\kappa$ with $\kappa$ the surface gravity. Here we work in the usual complexified set-up where the functions and their barred versions are complex and independent (e.g. $\lambda_{0}$ and $\bar{\lambda}_{0}$ are not complex conjugate). Furthermore, we will require the following constraints on the variations:
@@
[…]
@@
The second symplectic pair in (17) involves $\bar{p}^{A}$, whose explicit expression in terms of the spin coefficients will be provided in [47], and $\bar{m}^{0}_{A}$. As we will see in the next section, the latter canonical variable does not appear in the $Lw_{1+\infty}$ symmetry generators (24). Hence, it will not play any role in the computation of the $Lw_{1+\infty}$ charge algebra.
@@ V Canonical $Lw_{1+\infty}$ symmetries
The charges are labeled by $s=-1,0,1,\ldots$, the spin weight of their associated charge aspect $Q_{s}$. The $T_{s}$’s are the associated symmetry parameters. These are weighted scalars satisfying þ${\text{\th}}^{\prime}_{\mathscr{C}}T_{s}=0$, generalizing the time-independence condition of the parameters at null infinity. The global factor $r$ indicates that these charges are associated with the subleading phase space at the horizon (17). The correction terms in (19) are determined by requiring the criterium of conservation, $(i)$. The explicit form of these correction terms can be found in Appendix A up to spin $2$. Using the recursion relations (13), the commutation properties of the GHP operators (9) (see [58, 57]), and the self-duality conditions (14), we show that
@@
[…]
and their explicit expression appear in Appendix A up to spin $2$. The integral over $v$ converges for all $s$ if, as in [25, 26, 59, 28], we impose Schwartzian falloffs: $\lim_{v\to\pm\infty}\lambda_{0}\sim e^{-|v|^{2}}$. This is consistent with the action of the $Lw_{1+\infty}$ symmetries on the phase space if the parameters satisfy the wedge condition: ð$\eth_{\mathscr{C}}^{s+2}T_{s}=0$. Under this assumption, the soft terms in the fluxes (24) vanish and, in particular, $\mathcal{F}_{-1}=0$.
@@ Acknowledgements.
Appendix A Charges expressions
In this Appendix, we display the explicit expressions of the $Lw_{1+\infty}$ surface charges (19) up to spin $2$, but the procedure can be iterated for any spin $s$. In particular, these lower spin charges are enough to generate the whole tower of $Lw_{1+\infty}$ charges through the algebra (22).
$H_{-1}$ $=\frac{r}{8\pi G}\oint T_{-1}Q_{-1},$
$H_{0}$ $=\frac{r}{8\pi G}\oint T_{0}\left(Q_{0}-U\,\eth_{\mathscr{C}}Q_{-1}\right),$ (23)
$H_{1}$ $=\frac{r}{8\pi G}\oint T_{1}\Big{[}Q_{1}-U\eth_{\mathscr{C}}Q_{0}+\frac{U^{2}}{2}\eth_{\mathscr{C}}^{2}Q_{-1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2Q_{-1}({\text{\th}}_{\mathscr{C}}^{\prime})^{-1}\sigma_{0}\Big{]},$
$H_{2}$ $=\frac{r}{8\pi G}\oint T_{2}\Big{[}Q_{2}-U\eth_{\mathscr{C}}Q_{1}+\frac{U^{2}}{2}\eth_{\mathscr{C}}^{2}Q_{0}$
$\qquad\qquad-\frac{U^{3}}{6}\eth_{\mathscr{C}}^{3}Q_{-1}+3Q_{0}({\text{\th}}_{\mathscr{C}}^{\prime})^{-1}\sigma_{0}$
$\quad\quad-3\eth_{\mathscr{C}}Q_{-1}({\text{\th}}_{\mathscr{C}}^{\prime})^{-2}\sigma_{0}-2\eth_{\mathscr{C}}(Q_{-1}({\text{\th}}_{\mathscr{C}}^{\prime})^{-1}(U\sigma_{0}))\Big{]}.$
Here $U=\kappa^{-1}\left(-1+e^{-\kappa v}\right)=-v+\mathcal{O}(\kappa)$ is a “dressed time” at the horizon satisfying þ${\text{\th}}^{\prime}_{\mathscr{C}}U=1$, which generalizes the simple bare time condition $\partial_{v}v=1$. We also introduced (þ$({\text{\th}}^{\prime}_{\mathscr{C}})^{-1}$, the inverse operator of þ${\text{\th}}^{\prime}_{\mathscr{C}}$ satisfying þ${\text{\th}}^{\prime}_{\mathscr{C}}({\text{\th}}^{\prime}_{\mathscr{C}})^{-1}=1$.
The integrated fluxes defined in (21) read as
[…]
$\mathcal{F}_{1}$ $=\frac{r}{8\pi G}\int dv\oint T_{1}\Big{[}U\eth_{\mathscr{C}}(\sigma_{0}Q_{-2})+\frac{U^{2}}{2}\eth^{3}_{\mathscr{C}}Q_{-2}$
$\qquad\qquad\qquad\qquad\qquad\qquad+2\eth_{\mathscr{C}}Q_{-2}({\text{\th}}_{\mathscr{C}}^{\prime})^{-1}\sigma_{0}\Big{]},$
$\mathcal{F}_{2}$ $=\frac{r}{8\pi G}\int dv\oint T_{2}\Big{[}-\frac{U^{2}}{2}\eth_{\mathscr{C}}^{2}(\sigma_{0}Q_{-2})-\frac{U^{3}}{6}\eth_{\mathscr{C}}^{4}Q_{-2}$
$\qquad\qquad-3\sigma_{0}Q_{-2}({\text{\th}}_{\mathscr{C}}^{\prime})^{-1}\sigma_{0}-3\eth_{\mathscr{C}}^{2}Q_{-2}({\text{\th}}_{\mathscr{C}}^{\prime})^{-2}\sigma_{0}$
$\qquad\qquad\qquad\qquad\qquad-2\eth_{\mathscr{C}}(\eth_{\mathscr{C}}Q_{-2}({\text{\th}}_{\mathscr{C}}^{\prime})^{-1}(U\sigma_{0}))\Big{]}.$

### v2

@@ I Introduction
The goal of this Letter is to identify the celestial $Lw_{1+\infty}$ symmetries of black hole horizons. [^1] Most of the discussion in this paper applies to generic null hypersurfaces, not only to black hole horizons. For conciseness, we will slightly abuse the terminology of “horizons” throughout the text, as our ultimate goal is to apply this analysis to black hole horizons. More precisely, we establish two main results: $(i)$ We demonstrate a precise correspondence between the Ashtekar-Streubel phase space at null infinity and the subleading phase space near a null hypersurface, such as a black hole horizon. This is achieved using Penrose’s conformal compactification, the behavior of the Newman-Penrose (NP) formalism under Weyl rescalings, and covariant phase space techniques. $(ii)$ Leveraging this correspondence, we identify the $Lw_{1+\infty}$ charges at the horizon and show that, upon imposing self-duality conditions, their fluxes integrated over the horizon yield the canonical generators of the celestial symmetries. In particular, in the absence of radiation through the horizon, these charges are conserved, giving rise to an infinite tower of observables relevant for black hole physics.
@@ II NP formalism at null infinity
$m=\frac{1}{r_{\mathscr{I}}}m_{0}^{A}\partial_{A}+\mathcal{O}(r_{\mathscr{I}}^{-2}),\quad\bar{m}=\frac{1}{r_{\mathscr{I}}}\bar{m}_{0}^{A}\partial_{A}+\mathcal{O}(r_{\mathscr{I}}^{-2})$
@@
[…]
@@
    $\bm{\Omega}_{\mathscr{I}}=\frac{1}{8\pi G}\int_{\mathscr{I}}dvd^{2}x\sqrt{q}\delta\lambda_{1}\wedge\delta\sigma_{0}+c.c.$ (6)
@@ III From null infinity to the horizon
[…]
@@
constructed from the bare operators by correcting them with spin coefficients, playing the role of a Weyl connection, so that they have well-defined weights. In [62], we display the explicit expression of these correction terms.
We now have all the tools in hand to map the relevant part of the solution space at null infinity to the one at the horizon. We implement the Penrose conformal compactification [58, 53, 49, 57] on the (partially) off-shell radial expansions presented in Section II, as the Einstein equations themselves are not Weyl covariant. Starting from the radial expansion of the tetrad (2), and performing the compactification with conformal factor $\Omega$, see (7), we find
@@
$\sigma\to\sigma^{\prime}=\sigma_{0}+\mathcal{O}(r),\quad\lambda\to\lambda^{\prime}=\lambda_{0}+r\lambda_{1}+\mathcal{O}(r^{2}).$ (12)
@@
In the NP formalism, the tetrad and spin-coefficients equations are not Weyl covariant, but the reursion relations are. Indeed, using the conformal GHP operators (9), one can “covariantize” the recursion relations (5) to obtain
${\text{\th }}^{\prime}_{\mathscr{C}}Q_{s}=\eth_{\mathscr{C}}Q_{s-1}-(s+1)\sigma_{0}\,Q_{s-2}$ (13)
This covariantization is harmless at null infinity, as all the correction terms involved in this procedure vanish at $\mathscr{I}$. Exploiting the Weyl covariance of these expressions, we can perform the conformal rescaling and re-interpret these equations at the horizon. However, at finite distance, the correction terms in the GHP derivative operators with respect to the bare operators are crucial as the intrinsic geometry at the horizon is non-trivial. For example, one of these corrections involves the surface gravity of the black hole. Hence, (13) constitutes the recursion relations at the horizon and, in particular, $Q_{s}=\Psi_{2-s}^{0}$, $s=-2,\ldots,2$ are now the components of the Weyl tensor in the Taylor expansion (11), and the $Q_{s}$’s for $s>2$ appear in the expansion of $\Psi_{0}$. Notice that, for spin s=1,0,1,2s=-1,0,1,2, (13) are the Bianchi identities at finite distance.
@@ IV Subleading phase space
As mentioned in Section II, a key ingredient to identify the $Lw_{1+\infty}$ symmetries from a spacetime perspective at null infinity are self-duality conditions. These are necessary only at subleading order in the radial expansion of the spin coefficients and recursion relations, where the deviation between self-dual and full gravity starts to occur [32, 35]. At finite distance, self-duality conditions are already visible at leading order in $r$ since the intrinsic geometry on the null hypersurface contains genuine degrees of freedom involving both helicities. We propose conditions on the spin coefficients:
@@
where $\mu_{0}$ is the expansion of $\mathcal{H}$ and $\text{Re}(\gamma_{0})=-2\kappa$ with $\kappa$ the surface gravity. Here we work in the usual complexified set-up where the functions and their barred versions are complex and independent (e.g. $\lambda_{0}$ and $\bar{\lambda}_{0}$ are not complex conjugate). The conditions (14) ensure that ${\text{\th }}^{\prime}_{\mathscr{C}},\eth_{\mathscr{C}}]=0$. Furthermore, we will require the following constraints on the variations:
@@
$\bm{\Omega}=\frac{r}{8\pi G}\,\oint\int dv\,\delta\sigma_{0}\wedge\delta\lambda_{0}+\delta\bar{p}^{A}\wedge\delta\bar{m}^{0}_{A}+\mathcal{O}(r^{2})$ (17)
@@
The second symplectic pair in (17) involves $\bar{p}^{A}$, whose explicit expression in terms of the spin coefficients will be provided in [62], and $\bar{m}^{0}_{A}$. As we will see in the next section, the latter canonical variable does not appear in the $Lw_{1+\infty}$ symmetry generators. Hence, it will not play any role in the computation of the $Lw_{1+\infty}$ charge algebra.
@@ V Canonical $Lw_{1+\infty}$ symmetries
The charges are labeled by $s=-1,0,1,\ldots$, the spin weight of their associated charge aspect $Q_{s}$. The $T_{s}$’s are the associated symmetry parameters. These are weighted scalars satisfying ${\text{\th }}^{\prime}_{\mathscr{C}}T_{s}=0$, generalizing the time-independence condition of the parameters at null infinity. The global factor $r$ indicates that these charges are associated with the subleading phase space at the horizon (17). The correction terms in (19) are determined by requiring the criterium of conservation, $(i)$. The explicit form of these correction terms can be found in [62] up to spin $2$. Using the recursion relations (13), the commutation properties of the GHP operators (9) (see [36, 56]), and the self-duality conditions (14), we show that
@@
$\mathcal{F}_{s}=\int dv\partial_{v}H_{s}=H_{s}|_{v=+\infty}-H_{s}|_{v=-\infty}$ (21)
and their explicit expression appear in [62] up to spin $2$. The integral over $v$ converges for all $s$ if, as in [32, 35, 33, 45], we impose Schwartzian falloffs: $\lim_{v\to\pm\infty}\lambda_{0}\sim e^{-|v|^{2}}$. This is consistent with the action of the $Lw_{1+\infty}$ symmetries on the phase space if the parameters satisfy the wedge condition: $\eth_{\mathscr{C}}^{s+2}T_{s}=0$. Under this assumption, the soft terms in the fluxes vanish and, in particular, $\mathcal{F}_{-1}=0$.